OpenStudy (anonymous):

A person plays a game where he blindly picks a colored chip from a container. There are 20 chips in the container: 5 red, 11 green, and 4 blue. He gets $1 for a red chip, nothing for a green chip, and he has to pay $1.25 if he picks a blue chip. This is a fair game. true or false

OpenStudy (rushwr):

For the game to be fair it should have 50 50 winning and losing chances right? You get paid only if u pick a red one and choosing a red one out of all is having a \[\frac{ 5 }{ 20 }= \frac{ 1 }{ }\] chance right? So do u think it's fair ?

OpenStudy (rushwr):

\[\frac{ 1 }{ 4 } *\]

OpenStudy (anonymous):

thats 1/4 chance that doesnt seem fair

OpenStudy (rushwr):

yeah so it's not fair!